OpenStudy (anonymous):

find the derivative y=1/2ln(x)

OpenStudy (anonymous):

I got y'=1/x

OpenStudy (anonymous):

Then you did something wrong. Is the expression like this: \[y = \frac{ 1 }{ 2 }\ln(x)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 1/2x

OpenStudy (anonymous):

dy/dy=0

OpenStudy (anonymous):

Ok. Then you did it right, but forgot the constant \[\frac{ 1 }{ 2 }\]. Multiply with that and it should be right.

OpenStudy (anonymous):

So the answer is, as glamendo said, \[y' = \frac{ 1 }{ 2x }\]